please help me with this.... I really need help with this question
A company tests the heavy duty elastic bands it m,akes by measuring how much they strech when supporting various masses.This data shows the results of tests on one of the elastic bands.
mass (kg) 0 , 2 , 4 , 6 , 8
Length (cm) 6.2 , 9.6 , 13.0 , 16.4 , 19.8
a) Graph this relation
b)what does the point (0,6.2) represent?
c)find the slope of the grah.Is it constant?what does it represent?
d)write an equation for the length in terms of the Mass?
e)predict how long the elastic band will be when supporting a 10 kg mass?
f) If the length for an 8 kg mass were 19.0 cm, how would the answer to part e) change?

Answer by stanbon(75887)

A company tests the heavy duty elastic bands it m,akes by measuring how much they strech when supporting various masses.This data shows the results of tests on one of the elastic bands.
mass (kg) 0 , 2 , 4 , 6 , 8
Length (cm) 6.2 , 9.6 , 13.0 , 16.4 , 19.8
a) Graph this relation
This site allows for graphing functions but not points.
---------------------
b)what does the point (0,6.2) represent?
The elastic band is 6.2 cm long when it is not being stretched.
--------------------------------
c)find the slope of the graph.
When mass increases by 1 kg the length increases by 3.4/2 = 1.7
So the slope is 1.7
---------------
Is it constant?
Yes
--------------------
what does it represent?
As the weight increases by 1 kg the elastic band stretches 1.7 cm
-------------------------
d)write an equation for the length in terms of the Mass?
length = 1.7(weight)+6.2
--------------------------------
e)predict how long the elastic band will be when supporting a 10 kg mass?
length = 1.7*10+6.2
length = 17+6.2 = 23.2 cm
-------------------------------
f) If the length for an 8 kg mass were 19.0 cm, how would the answer to part e) change?
The slope would change if you use the points (0,6.2), (8,19)
It would be (19.2-6.2)/8 = 13/8
Then length = (13/8)weight+6.2
If weight = 10 kg, length = (13/8)*10+6.2 = 22.45
=================
Cheers,
Stan H.